OpenStudy (anonymous):

Any help is appreciated! 1/(x-1) - 2/x^2=0

OpenStudy (anonymous):

\[1\div(x-1)=2\div(x^2)\] it wl gv u d equsn=x^-2x+2=o use the formula -b+/-\[-b \pm \sqrt{b ^{2}-4ac di}\div2a\]

OpenStudy (anonymous):

i tried it but i didnt get an answer though

OpenStudy (anonymous):

No solution

jhonyy9 (jhonyy9):

you will get x*2 -2x +2 =0 --- -2 +/- (-4)*(1/2) 2(-1 +/- i ) x_1,_2 = ----------------- = ------------ = -1 +/- i 2 2